OpenStudy (anonymous):

Identify whether the series ∑[i=1,∞,8(5/6)]^(i-1) is a convergent or divergent geometric series and find the sum, if possible.

OpenStudy (anonymous):

i know its divergent series

terenzreignz (terenzreignz):

\[\sum_{n=1}^{\infty}8\left( \ \frac{5}{6} \right)^{n-1}\] is this it?

OpenStudy (anonymous):

yes @terenzreignz

terenzreignz (terenzreignz):

To be convergent, the sequence must converge to 0, does it?